OpenStudy (anonymous):

WILL MEDAL can someone explain to me how to graph linear equations when the equation is -x + 2y = 11

OpenStudy (anonymous):

@phi

OpenStudy (error1603):

-x + 2y = 11

OpenStudy (anonymous):

?

OpenStudy (mathmate):

There are many ways to graph an equation. When given in the above form, you can graph by plotting the x- y-intercepts. x+2y=11 The x-intercept is when y=0, so x+2(0)=11=>x=11, x-intercept is (11,0). The y-intercept is when x=0, so (0)=2y=11 means y=11/2, so y-intercept is (0,5.5). now plot the graph:  You can also convert the equation to slope-intercept form and then calculate the slope, and intercept before plotting, but more work (therefore more chances of error) in this case.

OpenStudy (anonymous):

it said it was wrong :(

OpenStudy (mathmate):

@chronic.haley Are you referring to the graph drawn above?

OpenStudy (anonymous):

no i did the problem exactly how they wanted it done and they said it was wrong lol

OpenStudy (anonymous):

it's okay tho, thanks

OpenStudy (mathmate):

What did you give for the answer?

OpenStudy (anonymous):

what is your ansers on a b c d